OpenStudy (arbyj):

find the slope of each line 14. y = -5 15. y = -5/2x - 5 16. x + y = -3 how do i do this?

OpenStudy (anonymous):

This dosent makes sense. You are missing something

OpenStudy (arbyj):

that's what the question says,

OpenStudy (anonymous):

the slope of y=-5 is 0. remember the general form is y=mx+b where m is the slope. in here m=0 and b = -5.

OpenStudy (anonymous):

another way to think about it - slope is how y changes when x changes. you see that here y is independent of x. and finally you can draw y=-5 and see that it is horizontal line - which is the same for every x.

OpenStudy (camerondoherty):

For the second one it is easier y = mx + b m = slope b = y-intercept so the equation is y = -5/2x - 5 -5/2 is in place of m, which is the slope So therefore -5/2 is the slope c: